### v1

# Knowledge Independence Breeds Disruption but Limits Recognition

Abstract
Recombinant growth theory highlights the pivotal role of cumulative knowledge in driving innovation  (?, ?, ?, ?). Although interconnected knowledge facilitates smoother dissemination  (?, ?), its connection to scientific disruption remains poorly understood. Here, we quantify knowledge dependence based on the degree to which references within a given paper’s bibliography cite one another. Analyzing 53.8 million papers spanning six decades, we observe that papers built on independent knowledge have decreased over time. However, propensity score matching and regression analyses reveal that such papers are associated with greater scientific disruption, as those who cite them are less likely to cite their references. Moreover, a team’s preference for independent knowledge amplifies its disruptive potential, regardless of team size, geographic distance, or collaboration freshness. Despite the disruptive nature, papers built on independent knowledge receive fewer citations and delayed recognition. Taken together, these findings fill a critical gap in our fundamental understanding of scientific innovation, revealing a universal law in peer recognition: Knowledge independence breeds disruption at the cost of impact.
With the shift in modern science toward large-scale, complex endeavors  (?, ?, ?), and the emergence of new scientific paradigms  (?), there is an increasing need to integrate diverse knowledge streams in scientific research  (?, ?). Recombinant growth theory  (?) posits that a larger body of knowledge provides more opportunities for novel combinations, thereby fostering greater innovation. However, empirical observations seem to contradict this prediction, as innovation has paradoxically slowed down even declined over time  (?, ?). This suggests that simply combining existing knowledge is insufficient to drive innovation, and that the way such knowledge is combined matters  (?).
Previous studies have analyzed the research strategies underpinning scientific papers, highlighting a widespread tendency to rely on well-established knowledge  (?, ?). This tendency extends to individual career trajectories, where scholars tend to exploit familiar topics rather than explore unfamiliar ones  (?, ?). Yet, innovative insights seem to emerge from unfamiliar topics, particularly through papers that draw on atypical combinations of knowledge  (?). Prior research has shown that novel combinations of research contents  (?), datasets  (?), methods  (?, ?), and references  (?, ?) are associated with greater scientific impact. Beyond conventional impact metrics, recent studies have linked such knowledge recombination to scientific disruption, i.e., the degree to which a paper overshadows the prior work it builds upon  (?, ?). Specifically, disruption has been associated with the novel combinations of references  (?) and research methods  (?).
Despite these numerous studies, one critical aspect of knowledge combination remains largely unexplored in the Science of Science literature to date—the notion of path dependence  (?). Intuitively, path dependence captures the idea that current and future decisions, outcomes, or developments in a system are shaped by the “path” (or sequence) of events or choices that preceded them  (?, ?, ?). In technical domains, this notion has been extensively studied, particularly in the context of patents. Research has shown that path dependence, the reliance on densely-interrelated patents when creating new ones, may foster incremental contributions by building on established technologies, reducing development costs, and promoting efficient progress  (?, ?, ?). However, it may also constrain innovation by locking inventors into existing trajectories, making it more challenging to explore novel or disruptive ideas outside established frameworks  (?, ?, ?). What remains unclear is whether this general law—that path dependence consolidates cumulative knowledge while path independence encourages innovations—extends from technology (i.e., patents) to science (i.e., papers).
Addressing the aforementioned gap would require a measure of path dependence that can be applied to the context of Science of Science. However, existing measures are predominantly qualitative, offering a dichotomous view in which paths are either dependent or independent, rather than a continuous perspective encompassing a spectrum of path dependence to independence  (?, ?, ?). The only continuous measures proposed in the literature operate at the meso-level, assessing the extent to which technical institutions rely on their past knowledge and experience  (?, ?). Yet, developing a continuous measure of path dependence or independence at the micro-level remains an open question.
Here, we introduce a formal definition of knowledge path independence in the Science of Science. Intuitively, this measure quantifies the extent to which the references in a paper are “independent” in the sense that they do not cite one another. For instance, if reference $A$ cites reference $B$, then this would be a knowledge path from $B$ to $A$. As such, knowledge path independence decreases as these paths grow longer and more frequent, signifying that the paper is built on a well-knitted body of knowledge. Throughout the remainder of this article, we will omit the term “path” and simply write “knowledge independence” to avoid potential confusion with the broader notion of paths in graph theory.
Armed with our measure of knowledge independence, we analyze 53.8 million publications from 1955 to 2017 in the Web of Science, and uncover a marked drop in knowledge independence across disciplines. At the same time, knowledge independence is becoming increasingly associated with scientific disruption. This relationship remains robust even when controlling for focal papers’ impact and novelty, their references’ age, impact, and disruption, as well as the composition of the team that creates the paper. Propensity score matching indicates a significant effect of knowledge independence on scientific disruption, while normalized regression analyses indicate that this effect is greater than those of the above factors. We further uncover the mediating role of knowledge independence in the relationship between disruption and team composition, regardless of the team size, geographic distance, or collaboration freshness. Finally, we find that papers built on independent knowledge receive fewer citations and delayed recognition, thereby explaining the negative association between disruption and impact. We confirm the robustness of all these findings when examining a different dataset focusing on publications, namely, SciSciNet  (?). Overall, our findings address a crucial gap in the fundamental understanding of scientific innovation, demonstrating that the recombination of independent knowledge breeds disruption, thereby becoming a knowledge broker that occupies a niche in the intellectual ecosystem. Yet, this very independence simultaneously increases the challenge of knowledge diffusion, leading to limited and delayed recognition.
Quantifying knowledge independence
To quantify knowledge independence in the context of bibliometrics, we develop a measure that relies on the citation network. This is because citations serve as tangible indicators of knowledge flow, reflecting the dependency of the citing paper on its references. Intuitively, for any given paper, the measure quantifies the degree to which its references are “independent” in the sense that they do not cite one another. To this end, let us introduce two types of references: An $i$-type reference is one that does not cite any other work within the same reference list, whereas a $j$-type reference is one that does cite at least one other work in this list. Then, knowledge independence (KI) is measured as the difference between the fraction of $i$-type and $j$-type references. More formally:
$\text{KI}=\frac{n_{i}-n_{j}}{n_{i}+n_{j}},$ (1)
where $n_{i}$ and $n_{j}$ are the numbers of $i$-type and $j$-type references, respectively. This yields a continuous value between $-1$ and $1$. On one end of the spectrum, we have papers in which none of the references cites any of the other references, corresponding to a KI value of $1$ (since $n_{j}=0$). On the other extreme, we have papers in which every reference (apart from the oldest one) cites at least one other reference, corresponding to a KI value that approaches $-1$ ($\frac{1-n_{j}}{1+n_{j}}$, since the oldest reference cannot cite any of the other ones, which yields $n_{i}=1$). Fig. 1A illustrates three examples, each depicting a focal paper (represented as blue diamond) along with its references (gray circles), demonstrating how KI changes with different values of $n_{i}$ and $n_{j}$. Throughout our analyses, we exclude papers that cite only one reference, since their KI is bound to be $1$ by definition, lacking any meaningful interpretation.
The quantification of KI makes it possible to systematically explore the distribution and evolution of knowledge independence in scientific papers. In general, the KI distribution seems to resemble a bell-shaped curve (Fig. 1B), punctuated by two outlier peaks at values 00 and $1$. This anomaly is due to the large number of papers with relatively short bibliographies (figure S1). For every such paper, the range of KI contains relatively few (if any) values other than 00 and $1$, e.g., for any paper with two references, $\text{KI}\in\{0,1\}$ (i.e., 00 and $1$ are the only possible values), and for any paper with four references, $\text{KI}\in\{-0.5,0,0.5,1\}$ (i.e., 00 and $1$ represent half the possible values). Additionally, $1$ is the only value that appears in the range of KI regardless of the number of references, while 00 is the only value (other than $1$) that appears in the range whenever the number of references is even. Fig. 1C depicts the evolution of KI. As can be seen, there has been a notable decline in papers’ KI over the past six decades, regardless of the discipline under consideration.
Knowledge independence is associated with scientific disruption
In principle, building on independent knowledge creates more opportunities for novel combinations, thereby fostering innovation. To test this hypothesis, we examined the correlation between KI and the disruption index  (?, ?). The disruption index quantifies the extent to which a paper introduces novel contributions that overshadow the prior work it builds upon, leading subsequent scholars to cite the paper itself without citing its references. This index belongs to $[-1,1]$, representing a spectrum of papers ranging from fully developmental ($-1$) to fully disruptive ($1$) (see the Disruption section in materials and methods).
KI is an individual measure of the extent to which a paper connects originally isolated knowledge whereas the disruption index collectively measures the peer scholars’ future appreciation of this paper. Seemingly, these two are not necessarily associated with each other. But our findings strongly support the hypothesis of innovation: papers with higher KI are more disruptive (Fig. 2A, blue). This relationship remains robust even when focusing exclusively on disruptive papers (disruption index $>0$), suggesting that increasing KI not only correlates with higher disruption scores but also raises the odds of a paper being disruptive (Fig. 2A, green). These results reveal the dual role of knowledge combinations: dependent knowledge enables incremental advancements within established bodies of research, while independent knowledge fosters innovative combinations that tend to disrupt existing scientific ideas. The pattern holds when controlling for the length of the reference list within the focal paper (figure S2), and when introducing numerous variations to the KI and disruption indices (definitions in materials and methods, and results in figure S3).
Next, we control for three properties of the focal paper: impact, publication date, scientific discipline. Starting with impact, it is measured as the number of citations the paper accumulates within the first five years post-publication, denoted by $C_{5}$. As impact increases, the association between disruption and KI becomes more pronounced, especially when the KI value is small (Fig. 2B). Moving on to publication date, we find that the relationship between KI and disruption has strengthened in recent decades (Fig. 2C). We also find a systematic decline in disruption over time, consistent with prior findings  (?). Next, we consider the discipline of the focal paper. As shown in Fig. 2D, the relationship between KI and disruption holds universally, from natural to social sciences, suggesting that it is a fundamental feature of the interplay between knowledge recombination and scientific disruption.
Having controlled for various properties of the focal paper, let us now control for the properties of the references within that paper. One aspect lies in the topological structure of the references, as reflected in the novelty index, capturing the degree of atypicality in the combinations of references at the journal level  (?) (details in materials and methods). The relationship between KI and disruption remains robust after controlling for novelty (figure S4B-C). Another aspect of the focal paper’s references lies in their statistical patterns, such as average age, average impact, and average disruption. In terms of reference age, both papers with high and low KI tend to cite relatively younger references compared to those cited by papers with intermediate KI values, i.e., those balancing dependent and independent knowledge (figure S5A). Furthermore, papers with higher KI exhibit a preference for less impactful (figure S6A), but more disruptive references (figure S7A). Importantly, the association between KI and disruption remains robust after controlling for reference age (figure S5B-C), reference impact (figure S6B-C), and reference disruption (figure S7B-C).
We further investigate whether team composition influences the relationship between KI and disruption. We measure team size  (?), geographic distance  (?), and collaborative freshness  (?) of the focal paper (definitions in materials and methods) and then control for these variables. Our results reveal that the positive correlation between disruption and KI persists across varying team size (figure S8). Similarly, the inherent advantages of onsite and fresh teams do not alter the positive correlation between KI and disruption (figure S9-S10). These findings well dispel the concerns that team composition might confound the observed relationship, affirming the robustness of the link between KI and disruption.
Motivated by the strong association between KI and disruption, we perform a matching experiment to estimate the effect size. To this end, we conduct randomized controlled trials via propensity score matching  (?) (PSM, details in materials and methods). In particular, each paper is assigned a propensity score based on a regression model accounting for covariates, allowing for the matching of papers in the treated and control groups with similar propensity scores. The effect size, characterized as the average treatment effect on the treated (ATT), is then estimated by comparing the disruption differences between matched pairs. Fig. 2E shows the pairwise ATTs across varying treated and control group sizes. The results reveal that for each controlled size, the ATT increases as the treated size grows, transiting from negative to positive when the treated size approaches the controlled size. This robustly confirms a positive effect of increasing KI on disruption.
The above findings further prompt us to ask whether KI plays a dominant role in driving disruption. To compare its influence with other variables, we employ normalized ordinary least-squares (OLS) regression (details in materials and methods) to model disruption as a function of KI alongside other variables. As summarized in Table 1, predictions from these models reveal that KI exerts a significant and dominant effect on disruption. While most other variables also show statistical significance, their impact is relatively minor. The key exception is the reference disruption, which exhibits an effect comparable to that of KI. Notably, as revealed that papers with higher KI tend to cite more disruptive references (figure S7A), suggesting that reference disruption may not be an independent driver of disruption but rather a consequence of KI. To investigate this further, we perform a mediation analysis  (?) (details in materials and methods). The results confirm that nearly half of the effect attributed to reference disruption is mediated by KI, while other variables show minimal mediating effect in this relationship compared to KI (table S1).
Small, onsite, and fresh teams prefer independent knowledge
Despite the extensive utilization and discussion on scientific disruption, one critical question remains unexplored: what type of papers tend to be disruptive? The revealed dominant effect of KI on disruption offers a quantitative answer, bridging the collective measure of peer recognition with an inherent property of the paper. This insight prompts us to check if previously identified correlates of disruption are, in fact, driven or mediated by KI.
The most well-known research on disruption stems from team science. Prior studies have highlighted the impact of team composition on disruption, emphasizing factors such as team size  (?), geographic distance  (?), and collaboration freshness  (?). These studies suggest that small, onsite, and fresh teams are more likely to produce disruptive research. Based on these insights and our discovery of the link between KI and disruption, we investigate whether such teams tend to build on independent knowledge. Indeed, Fig. 3A-C confirms that this has been the case for decades, indicating that KI may serve as a mediator in producing disruptive research. To explore this possibility, we analyze the relationship between disruption and team composition while controlling for KI. Our investigations (Fig. 3D-F) reveal that teams favoring higher KI systematically produce more disruptive research. However, once KI is controlled for, the slope of the disruption–team relationship becomes notably weaker compared to the uncontrolled case (inset, Fig. 3D-F), verifying the existence of the mediating role of KI. Mediation analyses further confirm the non-trivial mediating effect from KI, particularly when compared to other variables (table S2-S4). These findings suggest that a team’s preference for independent knowledge amplifies its disruptive potential, regardless of its size, geographic distance, or collaboration freshness.
Knowledge independence draws limited and delayed recognition
A well-documented phenomenon in the study of scientific impact is the negative correlation between a paper’s disruption index and its citation-based impact  (?, ?, ?). We verify that this pattern persists across journal ranks, publication years, and disciplines (Fig. 4A-C). However, the mechanism underlying this correlation remains to be explored. Since both the disruption and citation are the collective measure of the peer recognition, their correlation is hard to be untangled without linking them to the intrinsic property of the paper itself.
Given the advantages of interconnected knowledge in facilitating knowledge dissemination and adoption  (?, ?), it is reasonable to hypothesize that higher KI may hinder citation-based impact. To test this, for any given year $T$, we measure how KI is distributed across all papers published in that year, thereby forming a baseline distribution. Then, for each subsequent year, $T+t$ ($t\geq 1$), we examine the bibliographies of all papers published in that year, and extract only the references that were published in year $T$. Finally, we compute the KI distribution of the extracted references—where a paper may be counted multiple times if cited more than once—and compare it against the baseline distribution. The results show a systematic shift towards lower KI values, i.e., papers with negative KI values are consistently overrepresented among cited references compared to the baseline distribution, while those with positive KI values are consistently underrepresented (Fig. 4D). This pattern supports our hypothesis, indicating a negative correlation between KI and citation-based impact (Fig. 4E). In addition, as cited age $t$ increases, the KI distribution of yearly cited papers gradually converges towards the baseline distribution, suggesting that the citation disadvantage of higher-KI papers diminishes over time. This delayed recognition indicates that higher-KI papers are more likely to be “sleeping beauties”  (?), remaining dormant for longer periods before experiencing citation bursts (Fig. 4F).
The above analyses confirm that KI is associated with both limited and delayed academic recognition, suggesting that the negative relationship between disruption and impact may be driven by KI. To explore this possibility, we analyze the relationship between disruption and impact while controlling for KI. Our results (Fig. 4G) show that among papers with comparable impact, disruption systematically increases with KI. Notably, once KI is controlled for, the negative slope of the disruption-impact relationship becomes significantly weaker compared to the uncontrolled case (inset, Fig.4G), highlighting KI’s mediating role. Mediation analyses further confirm that KI accounts for a substantial portion of the observed negative relationship, exceeding the explanatory power of other covariates (table S5). These findings confirm our hypothesis that KI increases the challenge of knowledge diffusion while simultaneously enhancing the disruptive potential of scientific work.
Discussion
This paper offers new insights into a perplexing question in scientific innovation: Why has the explosion of knowledge been accompanied by a paradoxical decline  (?) in scientific innovation? Our analysis reveals a steady decline in research that builds on independent knowledge. Yet, such research occupies a favorable niche within the intellectual ecosystem, leveraging latent connections in prior knowledge to produce more disruptive outputs. But this comes at a cost: its overall impact remains limited compared to mainstream research that builds on interconnected knowledge. These findings reveal a universal law in scientific recognition: Knowledge independence breeds disruption but limits recognition.
Multiple studies have examined the relationship between the scientific innovation of a given paper and the topological structure of its bibliography. These studies suggest that the novelty index of the bibliography correlate with the paper’s disruption and impact  (?, ?). While our knowledge independence (KI) index also focuses on the paper’s bibliography, it differs substantially from the aforementioned index. More specifically, the novelty index assesses the atypicality of journal co-appearances within the bibliography, offering a coarse-grained characterization of meso-level relationships between combined knowledge. In contrast, our KI focuses on micro-level relationships by examining direct citation links between references, providing a more granular characterization of knowledge interconnection. The fundamental difference between these measures is reflected in their inverted U-shaped relationship (figure S4A).
Our work also complements prior studies examining the relationship between team composition and scientific disruption, which highlight the disruptive potential of small  (?), onsite  (?), and fresh teams  (?). In particular, we find that such teams exhibit a stronger preference to build on independent knowledge. Moreover, we demonstrate that the relationship between team composition and disruption weakens when controlling for KI. These findings reinforce our assertion that KI serves as a mediator, offering an additional boost to disruption. Consequently, our work suggests a practical strategy for scientists aiming to achieve disruptive discoveries: researchers should prioritize ideas composed of independent knowledge in their collaborative efforts, regardless of team size, location, or collaboration experience.
Finally, our work extends previous observations  (?, ?, ?) on the negative relationship between disruption and impact, demonstrating that this pattern persists across various attributes of papers, including journal rank, publication year, and discipline. But the negative relationship between disruption and impact diminishes when controlling for KI, highlighting the mediating role of knowledge independence in the paradoxical negative relationship between disruption and impact. A deeper insight into the dual role of knowledge independence in scientific innovation is that when a paper is grounded in independent knowledge, the non-redundant information it introduces allows it to function as a knowledge broker, attracting dedicated niche attention. However, the lack of explicit connections between the recombined knowledge simultaneously impedes its broader diffusion.
Although research agendas leveraging independent knowledge can foster scientific disruption and increase the likelihood of receiving prestigious awards (figure S22A), knowledge independence has notably declined across disciplines—from the natural to the social sciences—over the past six decades (Fig. 1C). One possible explanation behind this phenomenon could be the increasing pressure to “publish or perish”  (?, ?, ?). For example, career advancement—through promotions or funding—is often contingent on a steady publication record, which may disincentivize risk-taking in favor of safer, incremental research. We confirm this by observing that funding sources are more dedicated towards papers built on highly dependent knowledge (figure S22B). This funding-driven preference may create a reinforcing cycle that prioritizes well-established knowledge, making it more challenging for researchers to pursue disruptive ideas. Another possible explanation could be the increasing knowledge burden  (?). As the volume of scientific knowledge grows, scholars are increasingly driven to specialize  (?), which may limit their ability to traverse knowledge boundaries effectively  (?). To counteract these barriers, policymakers should consider implementing more flexible evaluation criteria that reward innovative, high-risk research beyond incremental contributions. Additionally, research service providers should develop advanced paper recommendation systems to help scholars discover and integrate independent knowledge more efficiently. Together, these initiatives can encourage research agendas that prioritize independent knowledge, ultimately fostering greater scientific disruption.
Figure 1: Quantifying knowledge independence. (A) Three examples of how knowledge independence (KI) is calculated for a given focal paper (blue diamond) based on its references (gray circles). A reference is of $i$-type (green) if it does not cite any other reference, or of $j$-type (orange) if it does. KI is then the difference in the fraction of $i$-type and $j$-type references. (B) The KI distribution of 29,267,547 papers with at least two references published between 1955 and 2017 in the Web of Science. (C) The evolving downtrend of KI over time across disciplines. Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure 2: Knowledge independence is associated with scientific disruption. (A) For the 27,769,300 papers in the Web of Science that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. (B) The association between KI and disruption persists, regardless of the impact of the focal paper. Here, impact is measured as the number of citations received within the first five years after publication, denoted by $C_{5}$. (C) The association between KI and disruption persists across decades. (D) The association between KI and disruption persists, regardless of discipline. Bootstrapped 95% confidence intervals are shown as shaded bands in panels A-D. (E) The ATT (average treatment effect on the treated) matrix of KI on disruption via propensity score matching (PSM). Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Within each controlled group, the ATT increases as the treated size grows, transitioning from negative to positive when the treated size approaches the controlled size. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Figure 3: Team’s preference for knowledge independence. (A-C), The KI percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Curves with larger bounds are displayed in the inset to improve visualization. Over time, the relative KI percentile decreases with team size (panel A) and team distance (panel B), and increases with team freshness (panel C); the only exception is observed for the 1950s in the team distance analysis (inset of panel B). (D-F), Solid lines depict the relationships between disruption and team properties with fixed KI values, alongside the uncontrolled cases (insets). Dashed lines represent linear fitted curves. Teams with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are flatter compared to the uncontrolled curves (insets). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure 4: Knowledge independence explains the negative relationship between disruption and impact. We observe a universal negative relationship between papers’ disruption and impact, regardless of the journal rank (panel A), time period (panel B), and discipline (panel C). Here, a paper’s impact is measured by the percentile of its total citations within the same publication year and discipline, and the journal rank is based on Scimago Journal Rank. (D) For papers published in the year $T=2000$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline (see figure S11 for different values of $T$). (E) The citation percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative citation percentile decreases with the KI percentile. (F) A paper with a higher KI is more likely to become a Sleeping Beauty (SB coefficient, blue curve, left y-axis) and takes a longer silence to usher in the citation burst (Awaking time, green curve, right y-axis). The SB coefficient and Awaking time  (?) capture a delayed scientific recognition by quantifying the convexity and the age of abrupt change in the distribution of annual citations of a paper. (G) Solid lines depict the relationships between disruption and impact with fixed KI values, alongside the uncontrolled case (inset). Dashed lines represent linear fitted curves. For papers of comparable impact, those with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are significantly flatter compared to the uncontrolled curve (inset). Bootstrapped 95% confidence intervals are shown as shaded bands.
Table 1: Normalized regression models of disruption and the influencing factors. We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption, while controlling for various confounding factors. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 Model 10 Model 11
KI 0.4333*** 0.4217*** 0.4332*** 0.4208*** 0.4312*** 0.4328*** 0.2910*** 0.4332*** 0.4317*** 0.4333*** 0.2768***
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
Log $C_{5}$0.0433*** -0.0318***
(0.000) (0.000)
Novelty 0.0057*** -0.0035
(0.000) (0.000)
Reference count -$0**$0205***
(0.000) (0.000)
Reference age 0.0577*** 0.0094***
(0.000) (0.000)
Reference $C_{5}$0.0205*** -0.0173***
(0.000) (0.000)
Reference disruption 0.2318*** 0.2265***
(0.000) (0.000)
Team size -$0**$0030***
(0.000) (0.000)
Team distance -0.0318*** -0.0231***
(0.000) (0.000)
Team freshness -$0**$0008**
(0.000) (0.000)
Fixed discipline Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Fixed year Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
N 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467 9857467
$R^{2}$ 0.223 0.224 0.223 0.224 0.226 0.223 0.247 0.223 0.224 0.223 0.249
Acknowledgments
We thank P. Holme and M.-Y. Luo for helpful discussions. We thank A.-L. Barabási for giving access to the Web of Science data. The support and resources of the High Performance Computing Center from Southwest University and New York University Abu Dhabi are gratefully acknowledged.
Funding:
The work is supported by National Natural Science Foundation of China (NSFC) (No.72374173), University Innovation Research Group of Chongqing (No.CXQT21005), Fundamental Research Funds for the Central Universities (No.SWU-XDJH202303), China Scholarship Council (CSC) program (No.202306990091), and Chongqing Graduate Research Innovation Project (No.CYB22129).
Author contributions:
X.-Y. Y., T. R. and T. J. conceived and designed the research. X.-Y. Y. analyzed the empirical data and performed the visualizations. X.-Y. Y. and T. J. validated the methods. X.-Y. Y., T. R. and T. J. wrote and edited the manuscript.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Our study draws on data from the Web of Science (WOS) and SciSciNet. The WOS dataset is not publicly available, and was used under license from their respective publishers. The SciSciNet dataset is accessible via the research  (?).
Supplementary materials
Materials and Methods
Figs. S1 to S22
Tables  S1 to S11
Supplementary Materials for
Knowledge Independence Breeds Disruption but Limits Recognition
Xiaoyao Yu, Talal Rahwan, Tao Jia
Corresponding author. Email: talal.rahwan@nyu.edu (T. R.); tjia@swu.edu.cn (T. J.)
This PDF file includes:
Materials and Methods
Figures  S1 to S22
Tables  S1 to S11
Materials and Methods
Data
Our data source is from the Web of Science (WOS), encompassing records from the Science Citation Index Expanded (SCIE), Social Sciences Citation Index (SSCI), and Arts & Humanities Citation Index (AHCI) databases. This dataset includes over 53.8 million publications from 1955 to 2017, with 34,353,967 articles, 1,968,358 letters, 1,934,079 proceedings papers, and 1,473,705 reviews. Other types of records that fall out of these “traditional” forms of articles are excluded from our analyses. To classify papers by academic discipline, we adopt the categorization scheme developed by Wu et al.  (?), which consolidates 258 subject categories into 14 major disciplines: Agriculture, Biology, Chemistry, Computer science, Medicine, Engineering, Environment sciences, Humanities, Law, Management, Mathematics, Multidiscipline, Physical sciences, and Social sciences. We also replicate our main findings on an additional dataset, SciSciNet  (?), which is developed from the Microsoft Academic Graph (MAG) dataset  (?). This dataset comprises approximately 134 million publications from 1950 to 2022, thereby complementing the WOS data with a larger data scale and more recent records. Similarly, we consider the “traditional” forms of articles in our analyses, such as journal and conference papers. The results based on SciSciNet are provided in the Supplementary Materials (figure S12-S22, table S6-S11).
Knowledge independence
Four alternative definitions of knowledge independence are proposed for comparison. KI is the principal definition used in the main text. $\text{KI}^{{}^{\prime}}$ simplifies the equation by only measuring the fraction of $i$-type references. $\text{KI}_{\rm{adj}}$ subtracts 1 from the count of $i$-type references, to account for the oldest reference in the bibliography, which cannot cite any other reference in that list. Finally, $\text{KI}^{{}^{\prime}}_{\rm{adj}}$ simplifies the previous equation by only measuring the fraction of $i$-type reference. More formally, the four alternative measures are defined as follows:
  • $\text{KI}=\dfrac{n_{i}-n_{j}}{n_{i}+n_{j}}$
  • $\text{KI}^{{}^{\prime}}=\dfrac{n_{i}}{n_{i}+n_{j}}$
  • $\text{KI}_{\rm{adj}}=\dfrac{(n_{i}-1)-n_{j}}{(n_{i}-1)+n_{j}}$
  • $\text{KI}^{{}^{\prime}}_{\rm{adj}}=\dfrac{n_{i}-1}{(n_{i}-1)+n_{j}}$
Disruption
The groundbreaking work of Funk and Owen-Smith  (?) proposed an index that analyzes patent citations to quantify technological change. This index was then advocated by the highly influential work of Wu et al.  (?), who coined the term “disruption index” and were the first to apply it in the analysis of scientific papers. The basic idea behind this index is that a disruptive work shifts the focus of subsequent research from its prior knowledge to itself. Therefore, the disruption index is calculated as $D=\dfrac{c_{i}-c_{j}}{c_{i}+c_{j}+c_{k}}$, where $c_{i}$ represents the number of works that only cite the focal work, $c_{j}$ represents the number of works that cite both the focal work and its references, and $c_{k}$ represents the number of subsequent works that cite the references of the focal work but not the focal work itself. There has been ongoing debate on the rationale behind the inclusion of the term $c_{k}$ in the formula  (?, ?, ?). Specifically, $c_{k}$ quantifies the attention directed towards the focal work’s references while bypassing the focal work itself. Under fixed $c_{i}$ and $c_{j}$, focal works with higher $c_{k}$ values would be considered less disruptive. However, a contradiction arises in cases of negative $D$, where an increase in $c_{k}$ paradoxically increases $D$ rather than decreasing it—contrary to the conceptual definition of disruption. To address this inconsistency, we adopt a modified definition of disruption that excludes the term $c_{k}$. This modified measure serves as the principal metric of disruption in our study, while the original definition is retained as an alternative for robustness check. Additionally, we implement an open time window to count the citations a focal work receives ($c^{o}$) and use a 5-year time window ($c^{5}$) as an alternative measure.
To summarize, the disruption measures that we consider are defined as follows: $D_{0}$ is the principal definition used in the main text. $D_{1}$ takes the form of original definition with an open time window $c^{o}$. $D_{2}$ is defined the same way as $D_{0}$, but with a different time window $c^{5}$. $D_{3}$ takes the form of original definition with a 5-year time window $c^{5}$. More formally, these measures are defined as follows:
  • $D_{0}=\dfrac{c^{o}_{i}-c^{o}_{j}}{c^{o}_{i}+c^{o}_{j}}$
  • $D_{1}=\dfrac{c^{o}_{i}-c^{o}_{j}}{c^{o}_{i}+c^{o}_{j}+c^{o}_{k}}$
  • $D_{2}=\dfrac{c^{5}_{i}-c^{5}_{j}}{c^{5}_{i}+c^{5}_{j}}$
  • $D_{3}=\dfrac{c^{5}_{i}-c^{5}_{j}}{c^{5}_{i}+c^{5}_{j}+c^{5}_{k}}$
Novelty
The seminal work of Uzzi et al. (?) demonstrated that papers built on atypical combinations of prior work are more likely to be highly influential. This finding is based on a novelty index, which quantifies the degree of atypicality in the combinations of references at the journal level. More specifically, this index is calculated by comparing the observed co-occurrence of journal pairs in actual citation networks with their expected co-occurrence under randomized citation networks. This comparison allows for computing the $Z$-score of each possible journal pair cited by a given focal paper. A high and positive $Z$-score indicates a conventional combination of journal pairs, while a low and negative $Z$-score reflects a novel combination that deviates significantly from expected co-occurrence patterns. Following the way that novelty is computed in the SciSciNet  (?) dateset, we calculate the $10_{\rm th}$ percentile of the $Z$-score distribution of the journal pairs cited by the focal paper, and apply its negative value to represent the $90_{\rm th}$ percentile of novelty. This approach underscores the relatively top atypical combinations of references within the paper, capturing its degree of novelty.
Team distance
Lin et al.  (?) reported the rise of remote collaborations that span locations and time zones. These collaborations are less likely to produce disruptive works as they tend to focus more on technical tasks rather than conceptual ones. The original calculation of team distance relies on the latitude and longitude of the affiliations listed in a paper’s byline. Due to the absence of coordinate information in the Web of Science dataset, we instead use the following hierarchical categorization of the affiliations:
  • 00: Sub-organization / Department level;
  • $1$: Organization / University level;
  • $2$: City level;
  • $3$: State / Province level;
  • $4$: National level;
  • $5$: International level.
Using this hierarchy, the distance of any given team is categorized based on the highest level within the team. For instance, if there are three authors, two of which are in the same university, while the third is in another university situated at the same city, the distance will be categorized as $2$, implying that this is a city-level collaboration.
Team freshness
Zeng et al.  (?) revealed a positive correlation between team freshness and scientific disruption, emphasizing the importance of new team dynamics in fostering innovative research. To quantify team freshness, we analyze the prior collaboration network of the authors involved in a given focal paper. In this network, the undirected edges represent prior collaborations between pairs of authors before coauthoring the focal paper. We classify team freshness into mutually exclusive categories based on the topological structure of the collaboration network, where $V$, $k$, and $n$ represent the node set, node degree, and network size, respectively. More specifically, the classification is as follows:
  • 00: All authors have previously collaborated with each other, $\{\forall v_{i}\in V,\ k_{i}=n-1\}$;
  • $1$: All authors have at least one prior collaboration link, $\{\forall v_{i}\in V,1\leq\ k_{i}\leq n-1\}$;
  • $2$: Some authors have at least one prior collaboration link, $\{\exists v_{i}\in V,\ k_{i}=0\}$;
  • $3$: No author has any prior collaboration link, $\{\forall v_{i}\in V,\ k_{i}=0\}$.
Propensity score matching (PSM)
Propensity score matching (PSM)  (?) has been extensively employed to estimate causal treatment effects in the Science of Science  (?, ?). It offers a robust approach to account for potential biases induced by confounding factors, particularly in scenarios where randomized experiments are impractical. In this study, we apply PSM to rigorously quantify the effect size of KI on scientific disruption. The procedure involves the following steps. We first categorize papers into the treated group and controlled group according to treatment size (KI). Subsequently, we estimate the propensity score for each paper by fitting a generalized linear model where the treatment size is the dependent variable. The model includes the following covariates: focal paper properties (impact $C_{5}$ and publication year), reference properties (novelty, reference count, average age, average impact $C_{5}$, and average disruption), and team properties (size, geographic distance, and collaboration freshness). For each paper $i$ in the controlled group, a counterpart from the treated group with a similar propensity score is matched. This matching ensures a balanced distribution of covariates across the two groups, effectively mimicking the conditions of a randomized controlled trial. After identifying $m$ matched pairs of papers, denoted as $M$, we calculate the Average Treatment Effect on the Treated as $\text{ATT}=\frac{1}{m}\sum_{i\in M}(Y^{1}_{i}-Y^{0}_{i})$, where $Y^{1}_{i}$ is the outcome (scientific disruption) for the treated paper and $Y^{0}_{i}$ is the outcome for the control paper. To enhance the robustness of PSM results, we perform multiple experiments for different pairs of treated and control groups. This produces a matrix of ATT values (Fig. 2E), providing a comprehensive and nuanced view of the effect size of KI on scientific disruption.
Normalized ordinary-least-squares (OLS) regression
To analyze the relationships between scientific disruption and various covariates, we employ normalized ordinary least squares (OLS) regression models. This approach allows us to evaluate whether KI has a dominant effect on disruption. The regression models include the following covariates: focal paper properties (impact $C_{5}$, discipline, and publication year), reference properties (novelty, reference count, average age, average impact $C_{5}$, and average disruption), as well as team properties (size, geographic distance, and collaboration freshness). All numerical variables are standardized by $Z$-score normalization to ensure comparability of effect coefficients across models. We start with a model that does not control for any covariates other than fixing discipline and publication year (S1). Then, we control for each of the following covariates separately (S2): focal paper’s impact ($C_{5}$), novelty, reference count, reference age, reference impact ($C_{5}$), reference disruption, team size, team distance, and team freshness. Finally, we consider a model that accounts for all those covariates (S3). The results are displayed in Table 1.
$D$ $=b_{0}+a_{0}\text{KI}+c_{0}F,$ (S1)
$D$ $=b_{0}+a_{0}\text{KI}+a_{i}C_{i}+c_{0}F,$ (S2)
$D$ $=b_{0}+a_{0}\text{KI}+\sum\nolimits_{i=1}^{9}a_{i}C_{i}+c_{0}F.$ (S3)
Mediation analysis
Over the past decades, mediation analysis has seen rapid methodological development  (?), becoming particularly prevalent in biomedical and social sciences. This approach is often applied to data from randomized controlled trials, aiming to investigate whether and how the effect of a target variable $X$ on an outcome $Y$ operates through a mediating variable $M$. Here we utilize the model-based mediation framework, comprising two parametric regression models: Outcome Model (S4) formulates the relation between the outcome $Y$ and the target variable $X$, as well as the mediator $M$, with covariates $C$ controlled. On the other hand, the Mediator Model (S5) formulates the relation between the mediator $M$ and the target variable $X$, also controlling for covariates $C$. The indirect effect (ACME) is defined as the product of the coefficient $\beta_{1}$ of $X$ in model (S5) and the coefficient $\theta_{2}$ of $M$ in model (S4). This represents the portion of $X$’s effect on $Y$ mediated through $M$. The direct effect (ADE) is defined as the coefficient $\theta_{1}$ of $X$ in model (S4), capturing the direct relationship between $X$ and $Y$, independent of $M$. The total effect (TE) is defined as the sum of the indirect effect (ACME) and direct effect (ADE). In all conducted mediation analyses, we standardize the numerical variables by $Z$-score normalization to ensure comparability of effect coefficients across models, with discipline and publication year controlled.
$Y$ $=\theta_{0}+\theta_{1}X+\theta_{2}M+\theta_{3}C,$ (S4)
$M$ $=\beta_{0}+\beta_{1}X+\beta_{2}C.$ (S5)
Figure S1: Distribution of reference list length. The distribution of reference count for 33,708,219 WOS articles follows a stretched exponential pattern, with a large number of papers containing relatively short reference lists.
Figure S2: The association between disruption and KI is robust across different reference list lengths. (A-B) When controlling for the percentile of reference count, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference count. (C-D) When controlling for the reference count within the bottom group ($<=10$), where each reference count induces a limited range of KI values, both the percentile and positive ratio of disruption continue to increase with KI across all groups of reference count.
Figure S3: The association between disruption and KI is robust across alternative measures. (A-D) Upper panel: from the perspective of percentile, all disruption measures (materials and methods) show a consistent increasing trend with all KI measures. Lower panel: from the perspective of positive ratio, all disruption measures similarly increase with all KI measures. The positive ratio represents the likelihood of having a higher proportion of $i$-type references compared to $j$-type references. Consequently, the variation curves for $D_{1}$ and $D_{3}$ align with those for $D_{0}$ and $D_{2}$, respectively.
Figure S4: The association between disruption and KI is robust when controlling for novelty. (A) The novelty index (materials and methods) within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on less novel combinations of references, whereas neutral papers—those maintaining a balance between dependent and independent knowledge—show a preference for more novel combinations of references. This pattern remains consistent across time periods. (B-C) When controlling for focal papers’ novelty, both the percentile and positive ratio of disruption continue to increase with KI across all levels of novelty. Additionally, papers relying on more novel combinations of references tend to exhibit slightly higher disruption.
Figure S5: The association between disruption and KI is robust when controlling for reference age. (A), We represent the reference age of a given focal paper by calculating the the average age of cited references. The reference age within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on relatively younger references, whereas neutral papers—those maintaining a balance between dependent and independent knowledge—show a preference for older references. This pattern remains consistent across time periods. (B-C), When controlling for reference age, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference age.
Figure S6: The association between disruption and KI is robust when controlling for reference impact. (A), We represent the reference impact of a given focal paper by calculating the average $C_{5}$ of cited references. The reference impact within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers that draw on more independent knowledge tend to cite less impactful references. This pattern remains consistent across time periods. (B-C), When controlling for reference impact, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference impact. Additionally, papers citing less impactful references tend to exhibit slightly higher disruption.
Figure S7: The association between disruption and KI is robust when controlling for reference disruption. (A) We represent the reference disruption of a given focal paper by calculating the average disruption of cited references. The reference disruption within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers that draw on more independent knowledge tend to cite more disruptive references. This pattern remains consistent across time periods. (B-C) When controlling for reference disruption, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference disruption. Additionally, papers citing more disruptive references tend to exhibit higher disruption.
Figure S8: The association between disruption and KI is robust when controlling for team size. (A-B) When controlling for team size, both the percentile and positive ratio of disruption continue to increase with KI across different team sizes. Additionally, papers produced by smaller teams tend to exhibit slightly higher disruption.
Figure S9: The association between disruption and KI is robust when controlling for team distance. (A-B) We categorize the team distance of a given focal paper by hierarchical categorization of the affiliations, where the value from 00 to $5$ refers to collaboration hierarchy from sub-organization level to international level (materials and methods). When controlling for team distance, both the percentile and positive ratio of disruption continue to increase with KI across different team distances. Additionally, papers produced by geographically closer teams tend to exhibit slightly higher disruption.
Figure S10: The association between disruption and KI is robust when controlling for team freshness. (A-B) We classify the team freshness of a given focal paper by the topological structure of the prior collaboration network, where the value from 00 to $3$ refers to network topology from a full-connected network to an isolated network (materials and methods). When controlling for team freshness, both the percentile and positive ratio of disruption continue to increase with KI across different team freshnesses. Additionally, papers produced by fresher teams tend to exhibit slightly higher disruption.
Figure S11: Papers with higher KI are underrepresented among cited references and receiving delayed citations. (A-F) For papers published in the year $T$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline.
Figure S12: The distribution and evolution pattern of knowledge independence in SciSciNet dataset. (A) The KI distribution of 49,579,961 papers with at least two references published between 1950 and 2021 in the SciSciNet. (B) The evolving downtrend of KI over time across disciplines (the only exception is observed for Art). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S13: Knowledge independence is associated with scientific disruption in SciSciNet dataset. (A) For the 42,468,785 papers in the SciSciNet that are cited at least once, the average disruption percentile (blue curve, left y-axis) and the disruption positive ratio (green curve, right y-axis) increase with the KI percentile. (B) The association between KI and disruption persists, regardless of the impact of the focal paper. Here, impact is measured as the number of citations received within the first five years after publication, denoted by $C_{5}$. (C) The association between KI and disruption persists across decades. (D) The association between KI and disruption persists, regardless of discipline. Bootstrapped 95% confidence intervals are shown as shaded bands in panels A-D. (E) The ATT (average treatment effect on the treated) matrix of KI on disruption via propensity score matching. Each controlled group is set as a baseline, and ATTs are calculated for comparisons between the baseline and each of the treated groups. Blue cells represent negative ATTs, while red ones represent positive ATTs, with color intensity proportional to the absolute value. Each ATT is tested against the null hypothesis (ATT equals 0) using a two-sided t-test. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Figure S14: Team’s preference for knowledge independence in SciSciNet dataset. (A-C) The KI values in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative KI percentile decreases with team size (panel A) and team distance (panel B), and increases with team freshness (panel C). Here we calculate the geographic team distance by the coordinate information of affiliations, and categorize teams into two types—onsite ($\leq 100$ km) and remote ($>100$ km) teams (panel B). The relationship between KI and average collaboration distance is displayed in inset of panel B. (D-F) Solid lines depict the relationships between disruption and team properties with fixed KI values, alongside the uncontrolled cases (insets). Dashed lines represent linear fitted curves. Teams with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are flatter compared to the uncontrolled curves (insets). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S15: Knowledge independence explains the negative relationship between disruption and impact in SciSciNet dataset. We observe a universal negative relationship between papers’ disruption and impact, regardless of the journal rank (panel A), time period (panel B), and discipline (panel C). Here, a paper’s impact is measured by the percentile of its total citations within the same publication year and discipline, and the journal rank is based on Scimago Journal Rank. (D) For papers published in the year $T=2000$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline (see figure S21 for different values of $T$). (E) The citation percentiles in each bin are rescaled by the average value of the respective time period to highlight the trends. Over time, the relative citation percentile decreases with KI percentile. (F) A paper with a higher KI is more likely to become a Sleeping Beauty (SB coefficient, blue curve, left y-axis) and takes a longer silence to usher in the citation burst (Awaking time, green curve, right y-axis). (G) Solid lines depict the relationships between disruption and impact with fixed KI values, alongside the uncontrolled case (inset). Dashed lines represent linear fitted curves. For papers of comparable impact, those with higher KI are consistently more disruptive. Moreover, the slopes of the fitted curves with fixed KI are significantly flatter compared to the uncontrolled curve (inset). Bootstrapped 95% confidence intervals are shown as shaded bands.
Figure S16: Distribution of reference list length in SciSciNet dataset. The distribution of reference count for articles recorded in SciSciNet follows a stretched exponential pattern, with a large number of papers containing relatively short reference lists.
Figure S17: The association between disruption and KI is robust across different reference list lengths in SciSciNet dataset. (A-B) When controlling for the percentile of reference count, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference count. (C-D) When controlling for the reference count within the bottom group ($<=10$), where each reference count induces a limited range of KI values, both the percentile and positive ratio of disruption continue to increase with KI across all groups of reference count.
Figure S18: The association between disruption and KI is robust across alternative measures in SciSciNet dataset. (A-D) Upper panel: from the perspective of percentile, all disruption measures show a consistent increasing trend with all KI measures. Lower panel: from the perspective of positive ratio, all disruption measures similarly increase with all KI measures. The positive ratio represents the likelihood of having a higher proportion of type $i$ references compared to type $j$ references. Consequently, the variation curves for $D_{1}$ and $D_{3}$ align with those for $D_{0}$ and $D_{2}$, respectively.
Figure S19: The association between disruption and KI is robust when controlling for focal paper’s novelty and reference’s properties in SciSciNet dataset. (A-D) When controlling for focal paper’s novelty (panel A), reference age (panel B), reference impact $C_{5}$ (panel C), and reference disruption (panel D), respectively, both the percentile (blue curves) and positive ratio (green curves) of disruption continue to increase with KI across all levels of above covariates.
Figure S20: The association between disruption and KI is robust when controlling for team compositions in SciSciNet dataset. (A-C) When controlling for team size (panel A), geographic distance (panel B), and collaboration freshness (panel C) separately, both the percentile (blue curves) and positive ratio (green curves) of disruption continue to increase with KI across all levels of above team compositions. Additionally, papers produced by small, onsite, and fresh teams tend to exhibit slightly higher disruption.
Figure S21: Papers with higher KI are underrepresented among cited references and receiving delayed citations in SciSciNet dataset. (A-G) For papers published in the year $T$, we measure the KI distribution among these papers as the baseline distribution (gray curve). Then, for all papers published in a subsequent year, $T+t$ ($t\geq 1$), we examine their bibliographies, extract the references that were published in $T$, and generate the KI distribution of those references (purple curve). Compared to the baseline, papers with negative KI are overrepresented while those with positive KI are underrepresented. Moreover, as the cited age $t$ increases, the corresponding distribution converges towards the baseline.
Figure S22: Knowledge independence is awarded but underfunded. (A) Based on the dataset for Nobel laureates  (?), We match 425 Nobel Prize-winning papers that cover the time period 1902–2016 in SciSciNet dataset, including 78 papers in Physics, 133 papers in Chemistry, and 214 papers in Medicine. The probability of winning the Nobel prize increases with the KI value of papers. (B) Based on the funding information recorded in SciSciNet dataset, we extract 263,5812 papers funded by NIH and 92,9177 papers funded by NSF. Here we repeat counting when a paper is funded by both agencies. The probability of being funded decreases with the KI value of papers.
Table S1: Normalized mediation regression on reference disruption. We perform normalized mediation analysis to assess the effect of reference disruption on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by reference disruption, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Reference disruption KI 0.2086*** 0.2318*** 0.4404***
(0.0006) (0.0009) (0.0007)
Reference disruption Novelty -0.0008*** 0.4412*** 0.4404***
(0.0000) (0.0007) (0.0007)
Reference disruption Reference count 0.0227*** 0.4177*** 0.4404***
(0.0005) (0.0009) (0.0007)
Reference disruption Reference age -0.0023*** 0.4427*** 0.4404***
(0.0002) (0.0007) (0.0007)
Reference disruption Reference $C_{5}$ 0.0005*** 0.4399 0.4404***
(0.0000) (0.0007) (0.0007)
Reference disruption Team size 0.0001*** 0.4403*** 0.4404***
(0.0000) (0.0007) (0.0007)
Reference disruption Team distance 0.0018*** 0.4386*** 0.4404***
(0.0000) (0.0007) (0.0007)
Reference disruption Team freshness -0.0000*** 0.4404*** 0.4404***
(0.0000) (0.0007) (0.0007)
Table S2: Normalized mediation regression on team size. We perform normalized mediation analysis to assess the effect of team size on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team size KI -0.0043*** -$0**$0123***
(0.0005) (0.0009) (0.0012)
Team size Novelty -$0**$0123*** -0.0123***
(0.0000) (0.0012) (0.0012)
Team size Reference count -0.0002*** -0.0121*** -0.0123***
(0.0001) (0.0012) (0.0012)
Team size Reference age -0.0018*** -0.0105*** -0.0123***
(0.0002) (0.0011) (0.0012)
Team size Reference $C_{5}$0.0004*** -0.0119 -0.0123***
(0.0000) (0.0012) (0.0012)
Team size Team distance -0.0037*** -0.0087*** -0.0123***
(0.0004) (0.0009) (0.0012)
Team size Team freshness -$0**$0123*** -0.0123***
(0.0000) (0.0012) (0.0012)
Table S3: Normalized mediation regression on team distance. We perform normalized mediation analysis to assess the effect of team distance on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team distance KI -0.0255*** -0.0318*** -0.0573***
(0.0003) (0.0006) (0.0007)
Team distance Novelty -$0**$0573*** -0.0573***
(0.0000) (0.0006) (0.0007)
Team distance Reference count -0.0084*** -0.0489*** -0.0573***
(0.0002) (0.0007) (0.0007)
Team distance Reference age -0.0027*** -0.0547*** -0.0573***
(0.0000) (0.0006) (0.0007)
Team distance Reference $C_{5}$0.0003*** -0.0571*** -0.0573***
(0.0000) (0.0007) (0.0007)
Team distance Team size -0.0007*** -0.0566*** -0.0573***
(0.0000) (0.0007) (0.0007)
Team distance Team freshness -$0**$0573*** -0.0573***
(0.0000) (0.0007) (0.0007)
Table S4: Normalized mediation regression on team freshness. We perform normalized mediation analysis to assess the effect of team freshness on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team distance, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team freshness, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team freshness KI 0.0007*** -$0**$0004***
(0.0003) (0.0006) (0.0006)
Team freshness Novelty $0**$0004*** -0.0004***
(0.0000) (0.0006) (0.0006)
Team freshness Reference count -0.0001*** -0.0003*** -0.0004***
(0.0001) (0.0006) (0.0006)
Team freshness Reference age -$0**$0003*** -0.0004***
(0.0000) (0.0006) (0.0006)
Team freshness Reference $C_{5}$$0**$0004*** -0.0004***
(0.0000) (0.0006) (0.0006)
Team freshness Team size -$0**$0004*** -0.0004***
(0.0000) (0.0006) (0.0006)
Team freshness Team distance -0.0003*** -0.0001*** -0.0004***
(0.0000) (0.0006) (0.0006)
Table S5: Normalized mediation regression on paper impact. We perform normalized mediation analysis to assess the effect of paper impact on disruption. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C5$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by paper impact, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Citation percentile KI -0.0989*** -$0**$0989***
(0.0003) (0.0006) (0.0006)
Citation percentile Novelty 0.0003*** -0.0991*** -0.0989***
(0.0000) (0.0006) (0.0006)
Citation percentile Reference count -0.0438*** -0.0551*** -0.0989***
(0.0009) (0.0011) (0.0006)
Citation percentile Reference age -0.0091*** -0.0897*** -0.0989***
(0.0001) (0.0007) (0.0006)
Citation percentile Reference $C_{5}$0.0007*** -0.0982*** -0.0989***
(0.0000) (0.0007) (0.0006)
Citation percentile Team size -0.0003*** -0.0986*** -0.0989***
(0.0000) (0.0006) (0.0006)
Citation percentile Team distance -0.0041*** -0.0948*** -0.0989***
(0.0001) (0.0007) (0.0006)
Citation percentile Team freshness $0**$0989*** -0.0989***
(0.0000) (0.0006) (0.0006)
Table S6: Normalized regression models of disruption and the influencing factors in SciSciNet dataset. We perform normalized ordinary-least-squares (OLS) regression analyses to examine the effect of KI on disruption in SciSciNet dataset, while controlling for various confounding factors. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6 Model 7 Model 8 Model 9 Model 10 Model 11
KI 0.3962*** 0.3831*** 0.3952*** 0.3766*** 0.3984*** 0.3960*** 0.2636*** 0.3959*** 0.3950*** 0.3925*** 0.2300***
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
Log $C_{5}$0.0469*** -0.0497***
(0.000) (0.000)
Novelty -0.0249*** -0.0232***
(0.000) (0.000)
Reference count -0.0623*** -0.0462***
(0.000) (0.000)
Reference age -0.0145*** -0.0416***
(0.000) (0.000)
Reference $C_{5}$0.0152*** -0.0218***
(0.000) (0.000)
Reference disruption 0.2237*** 0.2334***
(0.000) (0.000)
Team size -0.0096*** -0.0078***
(0.000) (0.000)
Team distance -0.0235*** -0.0149***
(0.000) (0.000)
Team freshness 0.0382*** 0.0329***
(0.000) (0.000)
Fixed discipline Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Fixed year Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
N 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656 19916656
$R^{2}$ 0.198 0.200 0.199 0.201 0.199 0.199 0.223 0.199 0.199 0.200 0.231
Table S7: Normalized mediation regression on reference disruption in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of reference disruption on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by reference disruption, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Reference disruption KI 0.1729*** 0.2237*** 0.3966***
(0.0004) (0.0006) (0.0005)
Reference disruption Novelty -0.0000*** 0.3966*** 0.3966***
(0.0000) (0.0005) (0.0005)
Reference disruption Reference count 0.0231*** 0.3735*** 0.3966***
(0.0001) (0.0005) (0.0005)
Reference disruption Reference age -0.0037*** 0.4003*** 0.3966***
(0.0001) (0.0005) (0.0005)
Reference disruption Reference $C_{5}$0.0003*** 0.3969*** 0.3966***
(0.0000) (0.0005) (0.0005)
Reference disruption Team size 0.0003*** 0.3963*** 0.3966***
(0.0001) (0.0005) (0.0005)
Reference disruption Team distance 0.0014*** 0.3952*** 0.3966***
(0.0000) (0.0005) (0.0005)
Reference disruption Team freshness 0.0037*** 0.3929*** 0.3966***
(0.0000) (0.0005) (0.0005)
Table S8: Normalized mediation regression on team size in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of team size on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team size, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team size KI -0.0106*** -0.0096*** -0.0203***
(0.0024) (0.0021) (0.0045)
Team size Novelty -0.0005*** -0.0198*** -0.0203***
(0.0001) (0.0044) (0.0045)
Team size Reference count -0.0007*** -0.0196*** -0.0203***
(0.0002) (0.0043) (0.0045)
Team size Reference age -0.0033*** -0.0169*** -0.0203***
(0.0007) (0.0038) (0.0045)
Team size Reference $C_{5}$0.0005*** -0.0198 -0.0203***
(0.0001) (0.0044) (0.0045)
Team size Team distance -0.0059*** -0.0144*** -0.0203***
(0.0012) (0.0033) (0.0045)
Team size Team freshness 0.0052*** -0.0255*** -0.0203***
(0.0012) (0.0057) (0.0045)
Table S9: Normalized mediation regression on team distance in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of team distance on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team distance, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team distance KI -0.0193*** -0.0235*** -0.0429***
(0.0002) (0.0004) (0.0004)
Team distance Novelty -0.0003*** -0.0425*** -0.0429***
(0.0000) (0.0004) (0.0004)
Team distance Reference count -$0**$0319*** -0.0429***
(0.0001) (0.0004) (0.0004)
Team distance Reference age -$0**$0419*** -0.0429***
(0.0000) (0.0004) (0.0004)
Team distance Reference $C_{5}$0.0002*** -0.0426*** -0.0429***
(0.0000) (0.0004) (0.0004)
Team distance Team size -$0**$0408*** -0.0429***
(0.0004) (0.0006) (0.0004)
Team distance Team freshness 0.0021*** -$0**$0429***
(0.0000) (0.0004) (0.0004)
Table S10: Normalized mediation regression on team freshness in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of team freshness on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C_{5}$, team size, and team distance, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by team freshness, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Team freshness KI 0.0342*** 0.0382*** 0.0725***
(0.0002) (0.0004) (0.0005)
Team freshness Novelty 0.0001*** 0.0724*** 0.0725***
(0.0000) (0.0005) (0.0005)
Team freshness Reference count 0.0097*** 0.0628*** 0.0725***
(0.0001) (0.0005) (0.0005)
Team freshness Reference age 0.0010*** 0.0715*** 0.0725***
(0.0000) (0.0005) (0.0005)
Team freshness Reference $C_{5}$0.0000*** 0.0725*** 0.0725***
(0.0000) (0.0005) (0.0005)
Team freshness Team size -0.0017*** 0.0742*** 0.0725***
(0.0004) (0.0006) (0.0005)
Team freshness Team distance -0.0013*** 0.0737*** 0.0725***
(0.0000) (0.0005) (0.0005)
Table S11: Normalized mediation regression on paper impact in SciSciNet dataset. We perform normalized mediation analysis to assess the effect of paper impact on disruption in SciSciNet dataset. The candidate mediating variables include: KI, novelty, reference count, reference age, reference $C5$, team size, team distance, and team freshness, with discipline and publication year controlled. ACME refers to average causal mediation effects driven from the mediating variable, ADE stands for average direct effects caused by paper impact, and TE is defined as the sum of the ACME and ADE. Each regression coefficient is tested against the null hypothesis (coefficient equals 0) using a two-sided t-test. Standard errors are provided in parentheses for each coefficient. Note that we do not apply adjustments for multiple hypothesis testing in this analysis. (*$p<0.05$, **$p<0.01$, ***$p<0.001$).
Independent variable Mediator ACME ($\beta_{1}\theta_{2}$) ADE ($\theta_{1}$) TE ($\theta_{1}+\beta_{1}\theta_{2}$)
Citation percentile KI -0.0873*** -0.0188*** -0.1061***
(0.0002) (0.0004) (0.0005)
Citation percentile Novelty -0.0019*** -0.1042*** -0.1061***
(0.0000) (0.0005) (0.0005)
Citation percentile Reference count -0.0456*** -0.0605*** -0.1061***
(0.0002) (0.0005) (0.0005)
Citation percentile Reference age -0.0036*** -0.1024*** -0.1061***
(0.0001) (0.0005) (0.0005)
Citation percentile Reference $C_{5}$0.0004*** -0.1057 -0.1061***
(0.0000) (0.0005) (0.0005)
Citation percentile Team size -0.0009*** -0.1052*** -0.1061***
(0.0002) (0.0005) (0.0005)
Citation percentile Team distance -0.0029*** -0.1032*** -0.1061***
(0.0000) (0.0000) (0.0005)
Citation percentile Team freshness -0.0038*** -0.1023*** -0.1061***
(0.0000) (0.0005) (0.0005)